OpenStudy (anonymous):

Solve: x + x/(x -2) = 2/(x - 2)

jimthompson5910 (jim_thompson5910):

x + x/(x -2) = 2/(x - 2) x(x -2) + (x -2)(x/(x -2)) = (x -2)(2/(x - 2)) ... Note: Multiply EVERY term by the LCD x-2 to clear out the fractions x(x -2) + x = 2 x^2 - 2x + x = 2 x^2 - 2x + x - 2 = 0 x^2 - x - 2 = 0 (x-2)(x+1) = 0 x-2 = 0 or x+1 = 0 x = 2 or x = -1 Now if x = 2, then the denominators in the original equation will be equal to zero. Remember that dividing by zero isn't defined. So x = 2 is not allowed. So the only solution is x = -1